answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

What is the probability that a ball on a roulette wheel will land on any number from 1 to 12?

There are 38 slots on a roulette wheel, two of them being "0" and "00", and the remainders being all numbers from one to thirty-six inclusive.

This means that your chances of getting any number from 1 to 12 are 12/38, or 6/19, which is is just a bit more than 31.5%

What is probability of throwing 9 heads in a row?

If you toss a coin often enough the probability is 1.

The probability of 9 H in the first 9 tosses is (1/2)9 = 1/512

What is probablitiy?

The chance something will happen

How many permutations are there for the letters in the word telephone?

The number of combinations of the letters TELEPHONE is 9 factorial, or 362,880. From that, however, we have to divide by 4 to account for the E being repeated twice, so the number of distinct combinations is 90,720.

What is the sample space when tossing 3 coins?

The sample space when tossing 3 coins is [HHH, HHT, HTH, HTT, THH, THT, TTH, TTT]

How is the term probability related to genetics?

The term probability is related to genetics because they both give guesses about how something that might be the outcome of something

What is the definition of probability in math?

Probability is the likelihood that some particular event will occur. It is expressed as a number between 0 and 1, where 0 means the event will not occur, and 1 means the event will occur. Values in between 0 and 1 indicate the relative likelihood, i.e. "probability", that the event will occur.

For instance, the probability of tossing heads on a fair coin toss is 0.5, the probability of rolling a 3 on a standard die is 1 in 6, or about 0.1667, and the probability of drawing an ace of spades from a standard deck of 52 cards is 1 in 52 or about 0.01923.

What is the probability that an event will definitely occur?

If the event will definitely occur, then its probability is 1.

Not asked, but answered for completeness sake - if the event will definitely not occur, then its probability is 0. All probabilities lie between 0 and 1, inclusive.

What is the probability of getting a license plate with a repeated letter or digit if you live in a state that has a 3 letters followed by 2 numbers?

The number of license plates without a repeated letter or digit is 26x25x24x10x9 = 1 404 000. The number of license plates without any restriction is 26x26x26x10x10 = 1 757 600.

Probability of a license plate without a repeated letter or digit = 1 404 000 /1 757 600 = 0.7988

Probability of a license plate with a repeated letter or digit = 1-0.7988 = 0.2022

Which has the best odds lotto Picking 6 out of 52 numbers or picking 5 out of 36 numbers?

The number of combinations of 52 numbers taken 6 at a time is 52 * 51 * 50 * 49 * 48 * 47 / 6 / 5 / 4 / 3 / 2 = 20,358,520. The number of combinations of 36 numbers taken 5 at a time is 36 * 35 * 34 * 33 *32 / 5 / 4 / 3 / 2 = 376,992.

The odds are better with 5 out of 36 numbers, but the payout is probably less as well.

What does it mean that the fan data are numerical and univariate?

It means that there is only one item of numerical information for each fan.

What does the z score of - 1.3 mean?

A z score of -1.3 means that the score is located at the negative 1.3 sigma level with respect to the mean.

Can you use this less than 5 percent rule for all normal probability distributions?

The answer depends on what "this less than 5 percent rule" is, in contrast to some other 5 percent rule!

What is the probability of getting a queen and a heart?

The probability of getting the queen of hearts is 1 in 52, or about 0.01923.

The probability of getting any queen is 4 in 52, or about 0.07692.

The probability of getting any heart is 13 in 52, or exactly 0.25.

Trending Questions
When A fair coin is tossed three times. Find the probability of getting a head on the first toss and tails on the second and third tosses? How do you make normal distribution curve in Excel? There are 2 red marbles 3 yellow marbles and 5 blue marbles in a bag determine the probability of drawing a yellow marble from the bag? What is the probability of getting prime number with a 6- sided dice? What is the probability of rolling a 6 an a fair sided cube? Which events in Napoleon's career occurred last in 1815? When a survey indicated that 65 of the family in a town own cars. find the probability that among 7 families selected at random in the town exactly 5 of them own cars? What is the role of an event official? A statement of probability that the experimental outcome was due to random variation is known as what? When preparing a PR and ampC you must? How do you derive variance of Pareto distribution? What does resolved mean? What is the probability of drawing a black king from a standard deck of 52 playing cards? How many possible outcomes does a set of 3 cards with different shapes on one side have? How may wild draw four cards in UNO? How do I use R to simulate a single fair coin toss My code needs to print Heads or Tails to the screen once every time it's run? Is areas within a continuous probability distribution represent probabilities? How many times would you expect draw the ace of spades if you have 13 cards in each of four suits and draw one card? What is the probability of landing on the number three? Can marbles be recycled?